# Samplet basis pursuit: Multiresolution scattered data approximation with sparsity constraints

### v3

@@ Abstract.
We consider scattered data approximation in samplet coordinates with $\ell_{1}$-regularization. The application of an $\ell_{1}$-regularization term enforces sparsity of the coefficients with respect to the samplet basis. Samplets are wavelet-type signed measures, which are tailored to scattered data. They provide similar properties as wavelets in terms of localization, multiresolution analysis, and data compression. By using the Riesz isometry, we embed samplets into reproducing kernel Hilbert spaces and discuss the properties of the resulting functions. We argue that the class of signals that are sparse with respect to the embedded samplet basis is considerably larger than the class of signals that are sparse with respect to the basis of kernel translates. Vice versa, every signal that is a linear combination of only a few kernel translates is sparse in samplet coordinates. Therefore, samplets enable the use of well-established multiresolution techniques on general scattered data sets.
@@ 1. Introduction
To enforce the sparsity, we apply regularization with respect to the $\ell^{1}$-norm, which improves the feature selection in comparison with the standard Tikhonov regularization. Dealing with such sparsity constraints in an efficient way is also mandatory for basis pursuit, that is, for decomposing given data into an optimal superposition of dictionary elements, where optimal means having the smallest $\ell^{1}$-norm of coefficients among all such decompositions, see [8, 27, 33] for instance. We demonstrate that basis pursuit problem can efficiently be solved by using a samplet basis. In particular, the samplet representation we obtain is extremely sparse.
In literature, the $\ell^{1}$-regularization term is treated numerically by using the iterative shrinkage-thresholding algorithm (ISTA), see [10], and its accelerated counterpart which is the fast iterative-shrinkage thresholding algorithm (FISTA), see [1]. Another approach to solve the underlying optimization problem is the alternating direction method of multipliers (ADMM), see [3], which we however do not consider here. Instead, we follow the idea of [15] and apply the semi-smooth Newton method, which corresponds to an active set method, cp. [9, 22]. As we will see, this leads to an efficient solver for the problem under consideration. Extensive numerical tests are presented and validate the feasibility and power of the proposed approach. Our benchmarks demonstrate the superiority of the multiresolution approach over the single-scale approach. As large-scale applications, we consider the surface reconstruction from noisy measurement data and the reconstruction of temperature data using a dictionary of multiple kernels.
@@ 2. Problem formulation
Depending on the choice of the kernel function, the linear system (6) is typically ill-conditioned and a suitable regularization is required to obtain a solution. Within this article, our focus is on $\ell^{1}$-regularization with respect to an appropriate multiresolution analysis. The corresponding multiresolution basis is introduced in the next section.
@@ 3.3. Characterization of the dual basis
$\psi_{j,k}$ $=\sum_{i}(\sigma_{j,k},\hat{k}_{i})_{\Omega}\hat{k}_{i}\in\mathcal{H}_{X},$
$\widetilde{\psi}_{j,k}$ $=\sum_{i}(\widetilde{\sigma}_{j,k},\hat{k}_{i})_{\Omega}\hat{k}_{i}\in\mathcal{H}_{X}$
@@ 4.1. Kernel ridge regression
$s_{h}({\boldsymbol{x}})=\sum_{i=1}^{N}\alpha_{i}\mathcal{K}({\boldsymbol{x}},{\boldsymbol{x}}_{i})=\sum_{i=1}^{N}\alpha_{i}\phi_{i}({\boldsymbol{x}})=\sum_{i=1}^{N}\beta_{i}\psi_{i}({\boldsymbol{x}}),$
@@ 4.2. Samplet basis pursuit
Note that the weight vector ${\boldsymbol{w}}\mathrel{\mathrel{\mathop{:}}=}[w_{i}]_{i}\in\mathbb{R}^{N}$ in (13) and (14), respectively, plays the role of the regularization parameter, where each coefficient is regularized individually. We refer to, e.g., [10, 25, 28] for the analysis of the regularizing properties and for appropriate parameter choice rules. Numerical algorithms to solve the optimization problems (13) and (14) are based on soft-thresholding. We recapitulate two efficient numerical realizations in the subsequent Section 5.
@@ 5.1. Fast iterative-shrinkage thresholding algorithm
       set ${\boldsymbol{\gamma}}_{1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\alpha}}_{0}$, ${\boldsymbol{\eta}}_{1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}{\boldsymbol{\gamma}}_{1}$, $t_{1}=1$ compute $\delta\mathrel{\mathrel{\mathop{:}}=}1/\big{\|}{\boldsymbol{K}}^{\Sigma}_{\varepsilon}\big{\|}_{2}^{2}$ for $i=1,2,\ldots$ do
             compute ${\boldsymbol{\alpha}}_{k}\mathrel{\mathrel{\mathop{:}}=}\operatorname{SS}_{\delta{\boldsymbol{w}}}\Big{(}{\boldsymbol{T}}^{\intercal}\big{(}{\boldsymbol{\eta}}_{k}-\delta({\boldsymbol{K}}^{\Sigma}_{\varepsilon})^{\intercal}({\boldsymbol{K}}^{\Sigma}_{\varepsilon}{\boldsymbol{\eta}}_{k}-{\boldsymbol{h}}^{\Sigma})\big{)}\Big{)}$ set $t_{k+1}\mathrel{\mathrel{\mathop{:}}=}\frac{1+\sqrt{1+4t_{k}^{2}}}{2}$ set ${\boldsymbol{\gamma}}_{k+1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\alpha}}_{k}+\frac{t_{k}-1}{t_{k+1}}({\boldsymbol{\alpha}}_{k}-{\boldsymbol{\alpha}}_{k-1})$ compute ${\boldsymbol{\eta}}_{k+1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}{\boldsymbol{\gamma}}_{k+1}$
@@ 5.2. Semi-smooth Newton method
       set ${\boldsymbol{\beta}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\beta}}_{0}}$ and ${\boldsymbol{r}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{F}({\boldsymbol{\beta}}_{0})$ while $\|{\boldsymbol{r}}\|>\textnormal{{tol}}$ do
@@
set ${\boldsymbol{M}}_{\mathcal{A}\mathcal{A}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{I}}_{\mathcal{A}}({\boldsymbol{K}}^{\Sigma}_{\varepsilon})^{\intercal}{\boldsymbol{K}}^{\Sigma}_{\varepsilon}{\boldsymbol{I}}_{\mathcal{A}}$ and ${\boldsymbol{M}}_{\mathcal{A}\mathcal{I}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{I}}_{\mathcal{A}}({\boldsymbol{K}}^{\Sigma}_{\varepsilon})^{\intercal}{\boldsymbol{K}}^{\Sigma}_{\varepsilon}{\boldsymbol{I}}_{\mathcal{I}}$ solve $\gamma{\boldsymbol{M}}_{\mathcal{A}\mathcal{A}}{\boldsymbol{\delta}}=\gamma{\boldsymbol{M}}_{\mathcal{A}\mathcal{I}}{\boldsymbol{r}}-{\boldsymbol{I}}_{\mathcal{A}}{\boldsymbol{r}}$ set ${\boldsymbol{\beta}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\beta}}+{\boldsymbol{\delta}}$ set ${\boldsymbol{r}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{F}({\boldsymbol{\beta}})$
@@
       set ${\boldsymbol{\beta}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\beta}}_{0}}$ and choose $\mu,\mu_{0}>1$ while $\mu>1$ do
             set ${\boldsymbol{\beta}}\mathrel{\mathrel{\mathop{:}}=}\texttt{MRSSN}\big{(}{\boldsymbol{K}}^{\Sigma}_{\varepsilon},{\boldsymbol{h}}^{\Sigma},\eta{\boldsymbol{w}},{\boldsymbol{\beta}},\gamma,\texttt{tol}\big{)}$ set $\mu\mathrel{\mathrel{\mathop{:}}=}\min\{1,\mu/\mu_{0}\}$
@@ 6.1. Benchmarks
Table 1. Results for sparse single-scale data.
Table 2. Results for sparse multiresolution data.
@@
Table 1. Results for sparse single-scale data.
Table 2. Results for sparse multiresolution data.
@@ 6.3. Sparse reconstruction of temperature data
The kernel $\mathcal{K}_{1}$ is a tensor product kernel comprised of a Matérn-3/2 kernel in space and a periodic kernel in time, both with relatively large correlation lengths. This kernel is intended to capture the smooth parts of the temperature distribution over time. The second kernel $\mathcal{K}_{2}$ is a very rough exponential kernel in space time and intended to capture sharp features. The overall time for the compression of both kernels is $508$s. The compressed matrix contains for $\mathcal{K}_{1}$ contains $507$ entries per row on average and the estimated relative compression error is $6.02\cdot{10^{-5}}$ in the Frobenius norm. The compressed matrix contains for $\mathcal{K}_{2}$ contains $939$ entries per row on average and the estimated relative compression error is $3.37\cdot{10^{-4}}$ in the Frobenius norm. We set $w_{i}=5\cdot 10^{-9}$ for $i=1,\ldots,N$ and MRSSN is stopped if $\|{\boldsymbol{r}}\|_{\infty}<9\cdot 10^{-7}$. The evaluation of the interpolant at all 12 458 880 points of the full data set approximately takes between $50$s and $73$s per time step using 32 cores.

### v4

@@ Abstract.
We consider scattered data approximation in samplet coordinates with $\ell_{1}$-regularization. The application of an $\ell_{1}$-regularization term enforces sparsity of the coefficients with respect to the samplet basis. Samplets are wavelet-type signed measures, which are tailored to scattered data. Therefore, samplets enable the use of well-established multiresolution techniques on general scattered data sets. They provide similar properties as wavelets in terms of localization, multiresolution analysis, and data compression. By using the Riesz isometry, we embed samplets into reproducing kernel Hilbert spaces and discuss the properties of the resulting functions. We argue that the class of signals that are sparse with respect to the embedded samplet basis is considerably larger than the class of signals that are sparse with respect to the basis of kernel translates. Vice versa, every signal that is a linear combination of only a few kernel translates is sparse in samplet coordinates.
@@ 1. Introduction
To enforce sparsity, we apply regularization with respect to the $\ell^{1}$-norm, which improves the feature selection in comparison with the standard Tikhonov regularization. Dealing with such sparsity constraints in an efficient way is also mandatory for basis pursuit, that is, for decomposing given data into an optimal superposition of dictionary elements, where optimal means having the smallest $\ell^{1}$-norm of coefficients among all such decompositions, see [8, 27, 33] for instance. We demonstrate that the basis pursuit problem can efficiently be solved by using a samplet basis. In particular, the samplet representation we obtain is extremely sparse.
In literature, the $\ell^{1}$-regularization term is treated numerically by using the iterative shrinkage-thresholding algorithm (ISTA), see [10], and its accelerated counterpart which is the fast iterative-shrinkage thresholding algorithm (FISTA), see [1]. Another approach to solve the underlying optimization problem is the alternating direction method of multipliers (ADMM), see [3], which we however do not consider here. Instead, we follow the idea of [15] and apply the semi-smooth Newton method, which corresponds to an active set method, cp. [9, 22]. As we will see, this leads to an efficient solver for the problem under consideration. Extensive numerical tests are presented and validate the feasibility and power of the proposed approach. Our benchmarks demonstrate the superiority of the multiresolution approach over the single-scale approach. As large-scale applications, we consider the surface reconstruction from measurements of the signed distance function and the reconstruction of temperature data using a dictionary of multiple kernels.
@@ 2. Problem formulation
Depending on the choice of the kernel function, the linear system (6) is typically ill-conditioned and a suitable regularization is required to obtain a solution. Within this article, our focus is on $\ell^{1}$-regularization with respect to an appropriate multiresolution analysis. The particular multiresolution basis is introduced in the next section.
@@ 3.3. Characterization of the dual basis
$\psi_{j,k}$ $=\sum_{i}(\hat{k}_{i},\sigma_{j,k})_{\Omega}\hat{k}_{i}\in\mathcal{H}_{X},$
$\widetilde{\psi}_{j,k}$ $=\sum_{i}(\hat{k}_{i},\widetilde{\sigma}_{j,k})_{\Omega}\hat{k}_{i}\in\mathcal{H}_{X}$
@@ 4.1. Kernel ridge regression
$s_{h}({\boldsymbol{x}})=\sum_{i=1}^{N}\alpha_{i}\mathcal{K}({\boldsymbol{x}}_{i},{\boldsymbol{x}})=\sum_{i=1}^{N}\alpha_{i}\phi_{i}({\boldsymbol{x}})=\sum_{i=1}^{N}\beta_{i}\psi_{i}({\boldsymbol{x}}),$
@@ 4.2. Samplet basis pursuit
Note that the weight vector ${\boldsymbol{w}}\mathrel{\mathrel{\mathop{:}}=}[w_{i}]_{i}$ in (13) and (14), respectively, plays the role of the regularization parameter, where each coefficient is regularized individually. We refer to, e.g., [10, 25, 28] for the analysis of the regularizing properties and for appropriate parameter choice rules. Numerical algorithms to solve the optimization problems (13) and (14) are based on soft-thresholding. We recapitulate two efficient numerical realizations in the subsequent Section 5.
@@ 5.1. Fast iterative-shrinkage thresholding algorithm
       set ${\boldsymbol{\gamma}}_{1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\alpha}}_{0}$, ${\boldsymbol{\eta}}_{1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}{\boldsymbol{\gamma}}_{1}$, $t_{1}=1$
      compute $\delta\mathrel{\mathrel{\mathop{:}}=}1/\big{\|}{\boldsymbol{K}}^{\Sigma}_{\varepsilon}\big{\|}_{2}^{2}$
      for $i=1,2,\ldots$ do
             compute ${\boldsymbol{\alpha}}_{k}\mathrel{\mathrel{\mathop{:}}=}\operatorname{SS}_{\delta{\boldsymbol{w}}}\Big{(}{\boldsymbol{T}}^{\intercal}\big{(}{\boldsymbol{\eta}}_{k}-\delta({\boldsymbol{K}}^{\Sigma}_{\varepsilon})^{\intercal}({\boldsymbol{K}}^{\Sigma}_{\varepsilon}{\boldsymbol{\eta}}_{k}-{\boldsymbol{h}}^{\Sigma})\big{)}\Big{)}$
            set $t_{k+1}\mathrel{\mathrel{\mathop{:}}=}\frac{1+\sqrt{1+4t_{k}^{2}}}{2}$
            set ${\boldsymbol{\gamma}}_{k+1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\alpha}}_{k}+\frac{t_{k}-1}{t_{k+1}}({\boldsymbol{\alpha}}_{k}-{\boldsymbol{\alpha}}_{k-1})$
            compute ${\boldsymbol{\eta}}_{k+1}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}{\boldsymbol{\gamma}}_{k+1}$
@@ 5.2. Semi-smooth Newton method
       set ${\boldsymbol{\beta}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\beta}}_{0}}$ and ${\boldsymbol{r}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{F}({\boldsymbol{\beta}}_{0})$
      while $\|{\boldsymbol{r}}\|>\textnormal{{tol}}$ do
@@
            set ${\boldsymbol{M}}_{\mathcal{A}\mathcal{A}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{I}}_{\mathcal{A}}({\boldsymbol{K}}^{\Sigma}_{\varepsilon})^{\intercal}{\boldsymbol{K}}^{\Sigma}_{\varepsilon}{\boldsymbol{I}}_{\mathcal{A}}$ and ${\boldsymbol{M}}_{\mathcal{A}\mathcal{I}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{I}}_{\mathcal{A}}({\boldsymbol{K}}^{\Sigma}_{\varepsilon})^{\intercal}{\boldsymbol{K}}^{\Sigma}_{\varepsilon}{\boldsymbol{I}}_{\mathcal{I}}$
            solve $\gamma{\boldsymbol{M}}_{\mathcal{A}\mathcal{A}}{\boldsymbol{\delta}}=\gamma{\boldsymbol{M}}_{\mathcal{A}\mathcal{I}}{\boldsymbol{r}}-{\boldsymbol{I}}_{\mathcal{A}}{\boldsymbol{r}}$
            set ${\boldsymbol{\beta}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\beta}}+{\boldsymbol{\delta}}$
            set ${\boldsymbol{r}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{F}({\boldsymbol{\beta}})$
@@
       set ${\boldsymbol{\beta}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{\beta}}_{0}}$ and choose $\mu,\mu_{0}>1$
      while $\mu>1$ do
             set ${\boldsymbol{\beta}}\mathrel{\mathrel{\mathop{:}}=}\texttt{MRSSN}\big{(}{\boldsymbol{K}}^{\Sigma}_{\varepsilon},{\boldsymbol{h}}^{\Sigma},\eta{\boldsymbol{w}},{\boldsymbol{\beta}},\gamma,\texttt{tol}\big{)}$
            set $\mu\mathrel{\mathrel{\mathop{:}}=}\min\{1,\mu/\mu_{0}\}$
@@ 6.1. Benchmarks
Table 1. Results for sparse single-scale data.
@@ 6.3. Sparse reconstruction of temperature data
The kernel $\mathcal{K}_{1}$ is a tensor product kernel comprised of a Matérn-3/2 kernel in space and a periodic kernel in time, both with relatively large correlation lengths. This kernel is intended to capture the smooth parts of the temperature distribution over time. The second kernel $\mathcal{K}_{2}$ is a very rough exponential kernel in space time and intended to capture sharp features. The overall time for the compression of both kernels is $508$s. The compressed matrix contains for $\mathcal{K}_{1}$ contains $507$ entries per row on average and the estimated relative compression error is $6.02\cdot{10^{-5}}$ in the Frobenius norm. The compressed matrix contains for $\mathcal{K}_{2}$ contains $939$ entries per row on average and the estimated relative compression error is $3.37\cdot{10^{-4}}$ in the Frobenius norm. We set $w_{i}=8\cdot 10^{-7}$ for $i=1,\ldots,N$ and MRSSN is stopped if $\|{\boldsymbol{r}}\|_{\infty}<9\cdot 10^{-7}$. The evaluation of the interpolant at all 12 458 880 points of the full data set approximately takes between $50$s and $73$s per time step using 32 cores.